How do u work the problem: log(x+3)+3log2=log16?
Answer by lwsshak3(6490)

How do u work the problem: log(x+3)+3log2=log16
**
log(x+3)+3log2=log16
log(x+3)+3log2-log16=0
place under single log
log[(x+3)(2^3)/16]=0
log[(x+3)(8)/16]=0
log[(x+3)/2]=0
convert to exponential form: base(10) raised to log of number(0)=number((x+3)/2)
10^0=(x+3)/2=1
x+3=2
x=-1